# Learning with Errors over Group Rings Constructed by Semi-direct ${}^{\dagger}$

### 01 Dec 2023 v2

@@ IV Hardness of decision $\mathsf{GR}\mbox{-}\mathsf{LWE}$
Let’s consider the group ring with an underlying group $G_{1}=\mathbb{Z}_{m}\ltimes\mathbb{Z}_{n}=\langle s\rangle\ltimes\langle t\rangle$, where $t$ has order $v$. Any element $\mathfrak{h}\in\mathbb{R}[G]$ of the form
@@
can also be regarded as a polynomial in $\mathbb{R}[x]$ with degree no greater than $v-1$ by replacing $t$ with the indeterminate $x$. It can be easily verified that there is at least one element in $\mathbb{R}[G]$ satisfying the evaluation (6) and (7) in Definition 12, which means $W_{r,\iota,T}$ is well-defined. In fact, we can choose the elements by the same method as Definition 11 in [30]. through Lagrange interpolation.
@@
Let $R=\mathbb{Z}[G]$ be a group ring of order $n$ with $G=\langle s\rangle\ltimes\langle t\rangle$ defined by (2), where $t$ is an element of order $v$. Let $\xi$ be a $v$-th primitive root of unity. Let $r(x)\in\mathbb{R}[x]$ be a polynomial with degree no greater than $v-1$, and let
$c=\left(\prod_{i=0}^{v-1}(r/\sqrt{v})(\xi^{i})\right)^{1/v}\geq 1.$
@@
and denote univariate polynomial $e_{j}(t):=e(\omega^{j},t)=f_{0}(t)+\omega^{j}f_{1}(t)+\cdots+\omega^{j(u-1)}f_{u-1}(t)$. Denote $\rho_{j}(e)=(e_{j}(0),e_{j}(\xi),e_{j}(\xi^{2}),\ldots,e_{j}(\xi^{v-1})):=(\rho_{j}^{(0)}(e),\rho_{j}^{(1)}(e),\ldots,\rho_{j}^{(v-1)}(e))$ for $j=0,1,\ldots,u-1$, where $\omega$ and $\xi$ are the $u$-th and $v$-th primitive roots of unity, respectively. In the following, we determine $\rho_{j}$ by determining its each coordinate $\rho_{j}^{(i)}$ for $i=0,1,2,\ldots,v-1$.
@@
First, we consider a fixed $j$, and then apply Proposition 1 to find a good enough approximation to $\frac{\rho_{j}^{(i)}(e)\pm\rho_{u-j}^{(i)}(e)}{2}$ for each $i$, which allows us to recover $e_{j}(t)\pm e_{u-j}(t)$ by solving a system of linear equations. Furthermore, given $e_{j}$ with respect to all $j$, we can recover $e$ efficiently except with negligible probability.
To achieve this goal, when $j=0,u/2$, we can use similar process of Lemma 9 in [30] to recover $\rho_{j}^{(i)}$ (In this case, $\rho_{j}^{(i)}$ is real). For the following discussion, we may assume $j\neq 0,u/2$. Define $k_{j}^{(i)}:\mathbb{C}\rightarrow\mathbb{R}[G_{1}]$ satisfying $\rho_{j}(k_{j}^{(i)}(z))=z\cdot\mathbf{e}_{i}+\bar{z}\cdot\mathbf{e}_{v-i}$, where $\mathbf{e}_{i}$ has 1 in the $i$-th coordinate and 0 otherwise, and we may restrict the image of $k_{j}^{(i)}$ within elements of $\mathbb{R}[G_{1}]$ which are of the form
@@
which has zero coefficients on $s^{i}t^{j}$ for any $11\leq i\leq u-$ 1 and $10\leq j\leq v-$ 1. On input $(z,m)$, the oracle $\mathcal{O}_{j}^{(i)}$ uses fresh Gaussian samples from $D_{\mathcal{I},\mathbf{A}_{k}^{(i)}}$, where $\mathbf{A}_{k}^{(i)}$ is the matrix representation of $r_{k}^{(i)}\in W_{r,\iota,T}$ and $(1+\iota)^{k}=\exp(m)$ as in Definition 12. Then it performs the transformation from Lemma 17 on these samples, the coset $\frac{e_{j}+e_{u-j}}{2}-\sum k_{j}^{(i)}(z_{j}^{(i)})+\mathcal{I}^{-1}$, parameter $r$ and matrix norm bound $d=\alpha q/(\sqrt{2}r)\cdot\omega(1)$, and convert them into $\mathsf{GR}\mbox{-}\mathsf{LWE}$ samples. Denote these samples by $A^{(i)}_{j,z,m}$. Then $\mathcal{O}_{j}^{(i)}$ calls the $R\mbox{-}\mathsf{DLWE}$ oracle on these samples and outputs 1 if and only if it accepts.
@@
Therefore, $\mathcal{O}_{j}^{(i)},\kappa,d^{\prime}$ correspond to a valid instance of $-(\exp(-\kappa),\exp(-\kappa),1+1/\kappa)\mbox{-}$OHCP with “hidden center” $\frac{\rho_{j}^{(i)}(e)+\rho_{u-j}^{(i)}(e)}{2}$, except with negligible probability.
@@
V Conclusion
@@ Appendix A Proof of Lemma 5
To determine all 1-dimensional irreducible representations of $G_{1}$, it suffices to consider the quotient group $G_{1}/[G_{1},G_{1}]$, where $[G_{1},G_{1}]$ deontes the commutator group of $G_{1}$, which is isomorphic to $\mathbb{Z}_{n}=\langle t\rangle$. Then $G_{1}/[G_{1},G_{1}]\cong\mathbb{Z}_{m}$ is also a cyclic group. According to the representation theory, the number of 1-dimensional representations of $G$ and $\mathbb{Z}_{m}$ are identical, i.e., $m=2r$. Furthermore, $\chi_{i}(s)$ is completely determined by all the (1-dimensional) representations of $\mathbb{Z}_{m}$, as stated in the lemma. Since $t$ and $st$ belong to the same conjugate class, then $\chi_{i}(t)=\chi_{i}(st)/\chi_{i}(s)=1$.
@@ Appendix B Proof of Lemma 8
Since $I$ is a right ideal of $\mathbb{Z}[G]$,

### 10 Mar 2026 v3

@@ IV Hardness of decision $\mathsf{GR}\mbox{-}\mathsf{LWE}$
Let’s consider the group ring with an underlying group of Type I: $G=\mathbb{Z}_{u}\ltimes\mathbb{Z}_{v}=\langle s\rangle\ltimes\langle t\rangle$, where $t$ has order $v$. Any element $\mathfrak{h}\in\mathbb{R}[G]$ of the form
@@
can also be regarded as a polynomial in $\mathbb{R}[x]$ with degree no greater than v1v-1 by replacing $t$ with the indeterminate $x$. It can be easily verified that there is at least one element in $\mathbb{R}[G]$ satisfying the evaluation (6) and (7) in Definition 12, which means $W_{r,\iota,T}$ is well-defined. In fact, we can choose the elements by the same method as Definition 11 in [30]. through Lagrange interpolation.
@@
Let $R=\mathbb{Z}[G]$ be a group ring of order $n$ with $G=\mathbb{Z}_{u}\ltimes\mathbb{Z}_{v}=\langle s\rangle\ltimes\langle t\rangle$ of Type I, where $t$ is an element of order $v$. Let $\xi$ be a $v$-th primitive root of unity. Let $r(x)\in\mathbb{R}[x]$ be a polynomial with degree no greater than v1v-1, and let
$c=\left(\prod_{i=0}^{v-1}(1/\sqrt{v})r(\xi^{i})\right)^{1/v}\geq 1.$
@@
and denote the univariate polynomial $e_{j}(t):=e(\omega^{j},t)=f_{0}(t)+\omega^{j}f_{1}(t)+\cdots+\omega^{j(u-1)}f_{u-1}(t)$. Denote $\rho_{j}(e)=(e_{j}(0),e_{j}(\xi),e_{j}(\xi^{2}),\ldots,e_{j}(\xi^{v-1})):=(\rho_{j}^{(0)}(e),\rho_{j}^{(1)}(e),\ldots,\rho_{j}^{(v-1)}(e))$ for $j=0,1,\ldots,u-1$, where $\omega$ and $\xi$ are the $u$-th and $v$-th primitive roots of unity, respectively. In the following, we determine $\rho_{j}$ by determining each of its coordinates $\rho_{j}^{(i)}$ for $i=0,1,2,\ldots,v-1$.
@@
Fix an index $j$ and apply Proposition 1 to obtain a sufficiently accurate approximation of $\frac{\rho_{j}^{(i)}(e)\pm\rho_{u-j}^{(i)}(e)}{2}$ for each $i$, which in turn enables recovery of $e_{j}(t)\pm e_{u-j}(t)$ via a system of linear equations. Having all $e_{j}$ for every $j$, we can efficiently reconstruct $e$ except with negligible probability.
First, we note that we can efficiently compute
$\frac{e_{j}(t)\pm e_{u-j}(t)}{2}+\mathcal{I}^{-1}$
from a representative of $e+\mathcal{I}^{-1}$ for all indices $j$. Consider a representative of the coset $e+\mathcal{I}^{-1}$, written as $e+x$ where $x\in\mathcal{I}^{-1}$. Let $x$ be written as a bivariate polynomial in the indeterminates $s$ and $t$, just as $e$ is: $x(s,t):=x_{0}(t)+sx_{1}(t)+\cdots+s^{u-1}x_{u-1}(t)$. As we defined $e_{j}(t)$ previously, we can define $x_{j}(t)$ in the same manner. For each $j$, we can compute $e_{j}(t)+x_{j}(t)$ by evaluating $e(s,t)+x(s,t)$ at $s=\omega^{j}$. We claim that $(e_{j}(t)+e_{u-j}(t)+x_{j}(t)+x_{u-j}(t))/2$ is a representative of the coset $\frac{e_{j}(t)+e_{u-j}(t)}{2}+\mathcal{I}^{-1}$, which directly achieves our goal. It remains to show that $(x_{j}(t)+x_{u-j}(t))/2\in\mathcal{I}^{-1}$. Since $x(s,t)\in\mathcal{I}^{-1}\supseteq\mathbb{Z}[G]$ and $x_{j}(t),x_{u-j}(t)$ have conjugate coefficients in $\mathbb{Z}[\omega]$, it follows that all the coefficients of $(x_{j}(t)+x_{u-j}(t))/2$ lie in $\mathbb{R}\cap\mathbb{Z}[\omega]=\mathbb{Z}$, which implies that $(x_{j}(t)+x_{u-j}(t))/2\in\mathbb{Z}[G]\subseteq\mathcal{I}^{-1}$. The same argument also applies to $\frac{e_{j}(t)-e_{u-j}(t)}{2}+\mathcal{I}^{-1}$.
To achieve our goal, when $j=0,u/2$, we can use similar process of Lemma 9 in [30] to recover $\rho_{j}^{(i)}$ (In this case, $\rho_{j}^{(i)}$ is real). For the following discussion, we may assume $j\neq 0,u/2$. Define $k_{j}^{(i)}:\mathbb{C}\rightarrow\mathbb{R}[G_{1}]$ satisfying $\rho_{j}(k_{j}^{(i)}(z))=z\cdot\mathbf{e}_{i}+\bar{z}\cdot\mathbf{e}_{v-i}$, where $\mathbf{e}_{i}$ has 1 in the $i$-th coordinate and 0 otherwise, and we may restrict the image of $k_{j}^{(i)}$ within elements of $\mathbb{R}[G_{1}]$ which are of the form
@@
which has zero coefficients on $s^{i}t^{j}$ for any $1\leq i\leq u-1$ and $0\leq j\leq v-1$. On input $(z,m)$, the oracle $\mathcal{O}_{j}^{(i)}$ uses fresh Gaussian samples from $D_{\mathcal{I},\mathbf{A}_{k}^{(i)}}$, where $\mathbf{A}_{k}^{(i)}$ is the matrix representation of $r_{k}^{(i)}\in W_{r,\iota,T}$ and $(1+\iota)^{k}=\exp(m)$ as in Definition 12. Then it performs the transformation from Lemma 17 on these samples, the coset $\frac{e_{j}+e_{u-j}}{2}-\sum k_{j}^{(i)}(z_{j}^{(i)})+\mathcal{I}^{-1}$, parameter $r$ and matrix norm bound $d=\alpha q/(\sqrt{2}r)\cdot\omega(1)$, and convert them into $\mathsf{GR}\mbox{-}\mathsf{LWE}$ samples. Denote these samples by $A^{(i)}_{j,z,m}$. Then $\mathcal{O}_{j}^{(i)}$ calls the $R\mbox{-}\mathsf{DLWE}$ oracle on these samples and outputs 1 if and only if it accepts.
@@
Therefore, $\mathcal{O}_{j}^{(i)},\kappa,d^{\prime}$ correspond to a valid instance of $(\exp(-\kappa),\exp(-\kappa),1+1/\kappa)\mbox{-}$OHCP with “hidden center” $\frac{\rho_{j}^{(i)}(e)+\rho_{u-j}^{(i)}(e)}{2}$, except with negligible probability.
@@
V Summary
V-A Relation to Module-$\mathsf{LWE}$
Similar to the discussion in [30], our construction is closely related to the Module-$\mathsf{LWE}$ problems. As an illustrative example, consider the construction arising from Type I groups. Let $R^{(1)}$ be the group ring associated with the Type I group $G_{1}=\mathbb{Z}_{m}\ltimes\mathbb{Z}_{n}=\langle s\rangle\ltimes\langle t\rangle$ specified in Section 3, and let $R_{q}^{(1)}=R^{(1)}/qR^{(1)}$. Define $\mathfrak{s}=\sum_{i=0}^{m-1}f_{i}(t)s^{i}$ and $\mathfrak{a}=\sum_{i=0}^{m-1}g_{i}(t)s^{i}$, where each $f_{i}(t),g_{i}(t)\ (0\leq i\leq m-1)$ is a polynomial over $\mathbb{Z}[t]/\langle t^{n/2}+1\rangle$ with coefficients sampled uniformly at random. Let $\mathfrak{b}=\sum_{i=0}^{m-1}b_{i}(t)s^{i}$ and let $\mathfrak{e}=\sum_{i=0}^{m-1}e_{i}(t)s^{i}$ be an error element in $R_{q}^{(1)}$. By the definition of the multiplication operation in the group ring, the equation $\mathfrak{b}=\mathfrak{s}\cdot\mathfrak{a}+\mathfrak{e}$ can can be written in matrix form:
$\left(\begin{array}[]{c}b_{0}\\ \vdots\\ b_{m-1}\end{array}\right)=\mathbf{M}\cdot\left(\begin{array}[]{c}g_{0}\\ \vdots\\ g_{m-1}\end{array}\right)+\left(\begin{array}[]{c}e_{0}\\ \vdots\\ e_{m-1}\end{array}\right),$
where each of the entries in $\mathbf{M}$ is a polynomial in $\mathbb{Z}[t]/\langle t^{n/2}+1\rangle$ that can be efficiently computed from $f_{0},\ldots,f_{m-1}$. Additionally, our construction can be viewed as a structured variant of the Module-$\mathsf{LWE}$ problem. This structure contrasts with the standard Module-$\mathsf{LWE}$ setting proposed in [9], in which the corresponding matrix $\mathbf{M}$ is typically assumed to be unstructured, with entries sampled independently and uniformly at random. In our case, the $m^{2}$ entries of $\mathbf{M}$ are fully determined by only $m$ polynomials. Consequently, to generate $O(mn)$ pseudorandom samples, our construction requires only $O(mn)$ random elements, namely, all the coefficients of $f_{0},\ldots,f_{m-1},g_{0},\ldots,g_{m-1}$, achieving a reduction by a factor of $m$ compared to the standard Module-$\mathsf{LWE}$ problem. The construction in [30] can be viewed as a special case of our framework when $m=2$, whereas our approach allows more flexible parameter choices. Moreover, according to the reduction result in [50], an instance of our rank-$m$ Module-$\mathsf{LWE}$ problem can be reduced to an instance of the Ring-$\mathsf{LWE}$ problem over a cyclotomic ring with modulus $q^{m}$.
V-B Conclusion
@@ Appendix A Proof of Lemma 5
To determine all 1-dimensional irreducible representations of $G_{1}$, it suffices to consider the quotient group $G_{1}/[G_{1},G_{1}]$, where $G_{1},G_{1}]$ deontes the commutator group of $G_{1}$, which is isomorphic to $\mathbb{Z}_{n}=\langle t\rangle$. Then $G_{1}/[G_{1},G_{1}]\cong\mathbb{Z}_{m}$ is also a cyclic group. According to the representation theory, the number of 1-dimensional representations of $G$ and $\mathbb{Z}_{m}$ are identical, i.e., $m=2r$. Furthermore, $\chi_{i}(s)$ is completely determined by all the (1-dimensional) representations of $\mathbb{Z}_{m}$, as stated in the lemma. Since $t$ and $st$ belong to the same conjugate class, then $\chi_{i}(t)=\chi_{i}(st)/\chi_{i}(s)=1$.
@@ Appendix B Proof of Lemma 8
Since II is a right ideal of $\mathbb{Z}[G]$,